A rectangle has width w. Its length is 1 more than twice its width. Find the perimeter of the rectangle in terms of w.

Found 2 solutions by algebrahouse.com, shree840:
Answer by algebrahouse.com(1659)

A rectangle has width w. Its length is 1 more than twice its width. Find the perimeter of the rectangle in terms of w.

w = width
2w + 1 = length {one more than twice width}
Perimeter = (2 x length) + (2 x width)
= 2(2w + 1) + 2w {two times length plus two times width equals perimeter}
= 4w + 2 + 2w {used distributive property}
= 6w + 2 {combined like terms}
www.algebrahouse.com

Answer by shree840(260)

length 2w+1
width w
perimeter is 2l+2w
=2(2w+1)+2w
=4w+2+2w
=6w+2